OpenStudy (anonymous):

A 52.0kg swimmer with an initial speed of 1.26m/s decides to coast until she comes to rest. If she slows with constant acceleration and stops after coasting 2.05m , what was the magnitude of the force exerted on her by the water?

OpenStudy (anonymous):

to find a: \[v_f^2 = v_i^2 + 2a \Delta s\] to find F: F = m*a

OpenStudy (anonymous):

because of Newton's 3rd where every action has opposite reaction, her force on the water = the force of the water on her

OpenStudy (anonymous):

So I have to find acceleration first? 0=1.26m/s+2a(delta)2.05?

OpenStudy (agent0smith):

delta s just means change in displacement... it shouldn't be in there once you plug in the displacement which was 2.05m.

OpenStudy (anonymous):

so \[a=\frac{ -1.26 }{ 2(2.05) }\]

OpenStudy (agent0smith):

Yep Then find F as Euler showed above.

OpenStudy (anonymous):

ok awesome thank you!

OpenStudy (anonymous):

So overall is the answer F=-15.98? The Mastering Physics says that answer is incorrect

OpenStudy (agent0smith):

You didn't square the velocity when finding a.

OpenStudy (anonymous):

Yes I just realized that!! I got F=20.3 and it is correct!

OpenStudy (agent0smith):

:)